OpenStudy (anonymous):

2 divide by r-2 subtrated from r+3 divide r-1

jhonyy9 (jhonyy9):

2 r+3 ----- - ------ = r-2 r-1 this is right ?

OpenStudy (anonymous):

the 2 is over r-2 - the r+3 is over the r-1

jhonyy9 (jhonyy9):

so than is right ,correct than this you need to calcule ,yes ?

jhonyy9 (jhonyy9):

so like a first step you need to calcule the denominator common for (r-2) and (r-1) how many will be ?

OpenStudy (anonymous):

(r-2)(r-1)

jhonyy9 (jhonyy9):

yes right so than what will be the next step ?

OpenStudy (anonymous):

2(r-1) divide by (r-2)(r-1) - (r+3)(r-2) divide by (r-2)(r-1)

jhonyy9 (jhonyy9):

yes is right but is more mathematicaly if you write it in this style : 2(r-1) (r+3)(r-2) --------- - ---------- = (r-2)(r-1) (r-2)(r-1) ok ? so the next step ?

OpenStudy (anonymous):

i got stuck after that stuck

jhonyy9 (jhonyy9):

comen please now this is very easy

jhonyy9 (jhonyy9):

2r-2-r^2 -2r +3r -6 ------------------- = (r-2)(r-1)

jhonyy9 (jhonyy9):

assume the same terms and than what you will get ?

jhonyy9 (jhonyy9):

courage please ,continue it

OpenStudy (anonymous):

2r-2-r^2+r-6

jhonyy9 (jhonyy9):

not is right 2r canceled out by -2r attention

OpenStudy (anonymous):

r *r is r^2

OpenStudy (anonymous):

r2+3r-8 / (r-2)(r-1)

jhonyy9 (jhonyy9):

so there is -r^2 not plus r^2 yes ?